OpenStudy (wolfgirl12):

what is the equation of the following graph? file:///C:/Users/candice/Documents/jasmine%20school/Capture.PNG

OpenStudy (aveline):

Please attach your image in another form ^^

OpenStudy (aveline):

Do you know how to find the equation of a parabola by using two points?

OpenStudy (wolfgirl12):

im pretty sure the equation of the parabola is (X +1)^2 = 8(Y +4) im just not sure how to get the second line part

OpenStudy (aveline):

The second line part? You mean the directrix?

OpenStudy (wolfgirl12):

ya

OpenStudy (wolfgirl12):

do i need to include that in as part of the graph or did i already figure it out?

OpenStudy (aveline):

Well, the directrix is closely related to the focus in the fact that it is the same distance from the focus. If you're just writing the equation, leave it out. If you're drawing a graph, make a dashed line at y=3 By the way, I'm not sure your equation is quite right.

OpenStudy (aveline):

*same distance from the vertex

OpenStudy (wolfgirl12):

oooh oops is it: (X +1)^2 = 4(Y +4) ?

OpenStudy (aveline):

Well, the vertex is at (1,4) (x-h)^2=4p(y-k) (x-1)^2=4(y-4)

OpenStudy (wolfgirl12):

how are you getting plus though? (the focus is 1,5) i just looked it up and it said it was subtraction not addition

OpenStudy (aveline):

Wait, what do you mean by getting plus?

OpenStudy (wolfgirl12):

oops sorry i meant to say minus.

OpenStudy (aveline):

Go to https://www.desmos.com/calculator This is the equation of the parabola: (x-1)^2=4(y-4) This is the focus: (1,5) And this is the directrix (y=3) Maybe looking at it visually will make most sense?

OpenStudy (wolfgirl12):

oh ya i guess you're right. thank you!

OpenStudy (aveline):

No problem :)